# Rule or Story, Which is a Better Commonsense Expression for Talking with Large Language Models?

### v1

@@ Abstract
[…]
@@ 1 Introduction
[…]
@@
However, commonsense is more than just rules Brachman and Levesque (2022). Humans acquire commonsense by recognizing prototypical patterns, extracting memories of similar past experiences, and contrasting them with the current novel situation to make decisions, as supported by psychological studies Schacter and Addis (2007); Klein (2004); Tulving (2002); Schank (1983); Schank and Abelson (1977). Our commonsense is often conveyed and passed down through stories such as myths and fairy tales Cassirer et al. (1946), with only a limited portion expressed in rules. Renowned AI theorist and cognitive psychologist Roger Schank argues in his book “Tell Me a Story: Narrative and Intelligence” that “knowledge is storiesSchank (1995). He emphasizes that humans struggle to learn and remember abstract rules derived from past experiences but can more easily remember a good story, because “stories give life to past experience”.
@@
To answer the questions, we systematically compare stories and rules as commonsense expressions for talking with LLMs. We use a total of 28 commonsense QA datasets for experiments. For the first question, we instruct LLMs to generate stories and rules based on commonsense questions, as shown in Figure 1. We compare the confidence and the accuracy of commonsense generation using stories and rules, showing that LLMs are more confident and more accurate at retrieving commonsense as stories than as rules. For the second question, we compare the confidence of generating the correct answers with stories or rules as contexts, showing that LLMs can more confidently leverage stories than rules for reasoning. The QA accuracy results further demonstrate that the story is a more effective commonsense expression for answering questions regarding daily events, while the rule is more effective for scientific commonsense QA. This phenomenon aligns with the reporting bias of commonsense in the text corpora. Moreover, stories and rules complement each other, i.e., combining them can further enhance the answer accuracy.
@@ 2 Background
There have been numerous works that inject commonsense rules into NLP models to improve commonsense reasoning, either by pre-training on knowledge bases Ma et al. (2021); Chang et al. (2020); Mitra et al. (2019); Zhong et al. (2019), or incorporating knowledge rules in the input of language models Shi et al. (2023); Wang et al. (2023c, f); Bian et al. (2021). There is also a chain of works that use graph-based reasoning for inference Wang et al. (2023e); Yasunaga et al. (2021); Lv et al. (2020); Lin et al. (2019). In this paper, we exploit the commonsense knowledge in the memory of LLMs as stories or rules to support commonsense reasoning.
Story generation by LLMs. Recent studies have found that LLMs can perform well on story generation Eldan and Li (2023); Wen et al. (2023); Peng et al. (2022). Bhandari and Brennan (2023) compare stories generated by OPT Zhang et al. (2022), LLaMA Touvron et al. (2023), and Alpaca Taori et al. (2023) with human-written stories, showing that these two kinds of stories exhibit a remarkable similarity in terms of readability and topics. The LLM-generated stories are even more accessible than traditional children’s stories. Xie et al. (2023) compared GPT-3 with story generation models before LLMs, demonstrating that LLMs generate stories of significantly high quality, even comparable with human authors. In this paper, we analyze and exploit the inherent commonsense embedded in the storytelling ability of LLMs.
@@ 3.2 Accuracy of Commonsense Generation
To assess which expression incorporates more accurate commonsense knowledge, we ask ChatGPT to determine if each story or rule aligns with commonsense, responding with either “yes” or “no” (prompt shown in Table 6 in Appendix A). We use the same 100 questions for each dataset in Section 3.1, randomly selecting one story and one rule for evaluation, which results in a total of 2,800 stories and 2,800 rules for each model.
@@ 4.1 Confidence in Commonsense Reasoning
${}^{\ddagger}$2Sci-elem 92.68 87.80 92.68 90.24 51.22 62.81 61.98 63.03 42.62 46.72 52.85 47.97
${}^{\dagger}$ellaSWAG-wikihow 73.62 69.45 62.75 68.29 28.89 31.44 30.63 28.69 26.43 25.78 25.64 25.42
@@
${}^{\ddagger}$ 75.70 77.43 77.14 79.14 27.09 40.53 45.08 41.64 22.59 28.57 26.14 27.68
${}^{\dagger\dagger}$omVE (Task A) 92.26 87.18 94.71 90.94 49.85 48.59 53.67 48.54 52.77 56.94 54.09 58.33
${}^{\ddagger}$-easy 92.46 92.11 93.86 92.81 59.30 63.15 68.28 65.29 51.67 53.68 54.74 56.59
@@ 6 Conclusion
The assessment of commonsense accuracy in Section 3.2 relies on automatic labeling by ChatGPT. Therefore, the accuracy presented in Table 1 serves solely for comparing stories and rules, and should not be regarded as absolute accuracy. To ensure the quality of this automatic evaluation, we conduct a manual review of the stories and rules labeled as incorrect by ChatGPT and confirm that they are indeed incorrect in terms of commonsense. However, due to its labor-intensive nature, manually labeling all generated stories and rules is impractical for us. Future studies need to incorporate human evaluation to provide a more comprehensive and nuanced understanding of the generated commonsense knowledge by LLMs.
@@ Appendix B Details of Commonsense QA Datasets
What is another name for a disk for storing information? A. computer store B. computer to store data C. computer hard drive D. cd player E. usb mouse
OpenBookQA
@@
Owls spend their nights A. tending to their homes B. sleeping in hollow logs C. scanning their territory for field mice D. hanging out with other owls
PIQA
@@
Where can I buy a tennis ball A. You can purchase a tennis ball at any sports store B. You can purchase a tennis racket at any sports store
SocialIQA
@@
Aubrey the officer pulled a driver over for speeding on the road. Why did Aubrey do this? A. find a safe place to pull the person over B. so people don’t drive to fast C. look up the person’s license plate number
ARC-easy
@@
Scientists at a local university have been studying the impact that people have on Earth. One of the areas being studied is how the burning of fossil fuels affects the environment. Which effect of fossil fuel burning have the scientists most likely evaluated? A. the production of nitrogen-fixing bacteria B. the mechanical weathering of roads C. the formation of acid rain D. the increase in runoff
ARC-challenge
@@
How should a line graph be used to display distance and time data for a moving object? A. The y-axis should be labeled as time, which is the dependent variable. B. The y-axis should be labeled as distance, which is the independent variable. C. The x-axis should be labeled as distance, which is the dependent variable. D. The x-axis should be labeled as time, which is the independent variable.
QASC
@@
What may renal failure be treated with? A. Laboratory B. Lymphocytes C. saves lives D. dialysis E. Lymph fluid F. dandelions G. ibuprofen H. Protein
AI2Sci-elem
@@
To make an electromagnet, a conductor should be coiled around - A. a glass tube B. an iron nail C. a roll of paper D. a wooden stick
AI2Sci-middle
@@
Which best describes the characteristics of a river basin? A. the land drained by a river and its tributaries B. the land formed when rivers create estuaries and marshes C. the land at the mouth of a river where water flows into the ocean D. the land formed as a result of a river flooding
WinoGrande
@@
She chose the black car over the green car, because the A. black car has more brighter color. B. green car has more brighter color.
WSC
@@
The user changed his password from "GrWQWu8JyC" to "willow-towered Canopy Huntertropic wrestles" as A. grwqwu8jyc was easy to remember. B. willow-towered canopy huntertropic wrestles was easy to remember.
NumerSense
a french horn has <how many> keys. (three)
HellaSWAG-act.net
@@
Another man practices hurling himself backward over a pole onto a gym mat inside of the gym. several more men A. practice hurling street hurling outside and on a gym floor. B. practice hurling while a coach’s hand watches. C. practice long jumps and backward jumps inside of the gym using the sandbox and gym mats as landing tools. D. practice pitches outside of the gym interior.
HellaSWAG-wikihow
@@
[header] How to do tiger eye hair [title] Purchase your hair dye. [step] Take a trip to your local drug store or beauty supply store. Depending on the look you’re going for, you may want to keep it simple and just choose one color, or you may want to buy four. A. Your hair will stick out more if you use a thicker dye, such as a mousse or gel. [title] Pour 3 ounces of red wine into your bowl. B. [substeps] In the tattoo artist’s shop you should find several different colors and strips, and apply those to your hair. Make sure the colors match what you want to do. C. It’s totally up to you! You can buy a blonde highlighting kit that will lighten pieces of your brown hair, auburn dye, gorgeous golden hues, soft brown dyes-whichever dye you think will look good in your hair. The darker your hair is, the less of an effect you’ll notice. D. [substeps] If you want to dye your hair yourself, make sure to use the formula before you apply the dye. Gels are sometimes recommended but are highly expensive, and can be difficult to find at grocery stores.
CommonsenseQA2.0
In the US a senator is a person elected to a six year term? A. yes B. no
SWAG
@@
Someone finds people playing chess at one of the long polished tables. She A. walks down the brightly decorated hall to join them. B. pats him in the gut with the box. C. faces the building, in wonder, someone and the other recruits stand around watching, uneasy. D. approaches two fat men wearing an earpiece into an office.
Com2Sense
@@
Because the drive was 20 miles long, Beth was able to make it to her destination in under 5 minutes. A. True B. False
SciQ
@@
What parts of a human possess the highest concentration of thermoreceptors? A. face and hair B. hand and ears C. face and ears D. hands and feet
QuaRel
@@
Lebron James a strong player for the Cavs battles Kevin Durant a thin player for a rebound. Who is likely to get the rebound? A. Durant B. Lebron
QuaRTz
@@
Long ago the surface of Venus warmed enough that greenhouse gases escaped into the atmosphere. As a result, the greenhouse effect on that planet A. increased B. decreased
CycIC
@@
Rob lauded Will. Charity chastised Will. Who made Will feel happy? A. Daisy B. Cliff C. Rob D. Charity E. Joy
ComVE (Task A)
@@
Which statement of the two is against common sense? A. The cleaner is in charge of the money at the store B. The cashier is in charge of the money at the store
COPA
@@
The woman was in a bad mood. What was the effect of this? A. She engaged in small talk with her friend. B. She told her friend to leave her alone.
PROST
@@
A person drops a bottle, a mirror, an egg, and a shirt from a balcony. Which object is the least likely to break? A. bottle B. mirror C. egg D. shirt
CODAH
Kieran is a whale. Kieran A. is a mammal B. is a dog C. has six human kids D. is a orange
SCT
@@
I wanted to buy a video game console. I asked my parents, and they came up with an idea. They said if I did my chores, I would be given money to save. I did my chores without being asked every week for a whole summer. What is the end of this story? A. My parents gave me enough money to buy the console. B. At the end of the summer I gave the money back to my parents.
$\alpha$NLI
@@
The beginning of the story: Jim got ready for his first date.The ending of the story: Since then, she has ignored all of Jim’s text messages. What happened between the begining and the end of the story? A. Jim’s date wasn’t attracted to him. B. Jim went on the date and said he didn’t like the girl.
WinoVenti
The walnut was painted. The walnut is A. edible B. toxic
@@ Appendix C Further Analyses of Generated Stories
C.1 Accuracy Differences between Using Stories and Rules for ChatGPT and Alpaca

### v2

@@ Abstract
Ning Bian1,2, Xianpei Han1,2,thanks: Corresponding Author, Hongyu Lin2, Yaojie Lu2, Ben He1,2, Le Sun1,2 1University of Chinese Academy of Sciences, Beijing, China 2Institute of Software, Chinese Academy of Sciences, Beijing, China bianning21@mails.ucas.ac.cn {xianpei,hongyu,yaojie,sunle}@iscas.ac.cn  benhe@ucas.ac.cn
@@ 1 Introduction
Figure 1: Comparison between rules and a story written by ChatGPT. The rules only provide useful knowledge until the 4th rule and also include an incorrect answer option, “classroom”. The story presents a detailed scenario where an adult uses glue sticks in an office.
@@
However, commonsense is more than just rules Brachman and Levesque (2022); Mostafazadeh et al. (2020). Humans acquire commonsense by recognizing prototypical patterns, extracting memories of similar past experiences, and contrasting them with the current novel situation to make decisions, as supported by psychological studies Schacter and Addis (2007); Klein (2004); Tulving (2002); Schank (1983); Schank and Abelson (1977). Our commonsense is often conveyed and passed down through stories such as myths and fairy tales Cassirer et al. (1946), with only a limited portion expressed in rules. Renowned AI theorist and cognitive psychologist Roger Schank argues in his book “Tell Me a Story: Narrative and Intelligence” that “knowledge is storiesSchank (1995). He emphasizes that humans struggle to learn and remember abstract rules derived from past experiences but can more easily remember a good story, because “stories give life to past experience”.
@@
To answer the questions, we systematically compare stories and rules as commonsense expressions for talking with LLMs. We use a total of 28 commonsense QA datasets for experiments. For the first question, we instruct LLMs to generate stories and rules based on commonsense questions, as shown in Figure 1. We compare the confidence and the accuracy of commonsense generation using stories and rules, showing that LLMs are more confident and more accurate at retrieving commonsense as stories than as rules. For the second question, we compare the confidence of generating the correct answers with stories or rules as contexts, showing that LLMs can more confidently leverage stories than rules for reasoning. The QA accuracy results further demonstrate that the story is a more effective commonsense expression for answering questions regarding daily events, while the rule is more effective for scientific commonsense QA. This phenomenon aligns with the reporting bias of commonsense in the text corpora. Moreover, stories and rules can complement each other, i.e., combining them can further enhance the answer accuracy.
@@ 2 Background
There have been numerous works that inject commonsense rules into NLP models to improve commonsense reasoning, either by pre-training on knowledge bases Ma et al. (2021); Chang et al. (2020); Mitra et al. (2019); Zhong et al. (2019), or incorporating knowledge rules in the input of language models Shi et al. (2023); Wang et al. (2023c, f); Lal et al. (2022); Bian et al. (2021). There is also a chain of works that use graph-based reasoning for inference Wang et al. (2023e); Yasunaga et al. (2021); Lv et al. (2020); Lin et al. (2019). In this paper, we exploit the commonsense knowledge in the memory of LLMs as stories or rules to support commonsense reasoning.
Story generation by LLMs. Generating fluent and coherent stories using NLP models has been a crucial component of computational creativity. Recent studies have found that LLMs can perform well on story generation Eldan and Li (2023); Wen et al. (2023); Peng et al. (2022). Bhandari and Brennan (2023) compare stories generated by OPT Zhang et al. (2022), LLaMA Touvron et al. (2023), and Alpaca Taori et al. (2023) with human-written stories, showing that these two kinds of stories exhibit a remarkable similarity in terms of readability and topics. The LLM-generated stories are even more accessible than traditional children’s stories. Xie et al. (2023) compared GPT-3 with story generation models before LLMs, demonstrating that LLMs generate stories of significantly high quality, even comparable with human authors. In this paper, we analyze and exploit the inherent commonsense embedded in the storytelling ability of LLMs.
@@ 3.2 Accuracy of Commonsense Generation
To assess which expression incorporates more accurate commonsense knowledge, we ask ChatGPT to determine if each story or rule aligns with commonsense, responding with either “yes” or “no” (prompt shown in Table 6 in Appendix A). We use the same 100 questions for each dataset in Section 3.1, randomly selecting one story and one rule for evaluation, which results in a total of 2,800 stories and 2,800 rules for each model. It is important to note that these annotations are used solely for the purpose of comparing stories and rules regarding commonsense accuracy under the same conditions, both annotated by ChatGPT, rather than implying absolute accuracy. Further human evaluation results for a small set of stories and rules are shown in Table 9 in Appendix C.1.
@@ 4.1 Confidence in Commonsense Reasoning
AI2Sci-elem 92.68 87.80 92.68 90.24 51.22 62.81 61.98 63.03 42.62 46.72 52.85 47.97
HellaSWAG-wikihow 73.62 69.45 62.75 68.29 28.89 31.44 30.63 28.69 26.43 25.78 25.64 25.42
@@
QASC 75.70 77.43 77.14 79.14 27.09 40.53 45.08 41.64 22.59 28.57 26.14 27.68
††ComVE (Task A) 92.26 87.18 94.71 90.94 49.85 48.59 53.67 48.54 52.77 56.94 54.09 58.33
ARC-easy 92.46 92.11 93.86 92.81 59.30 63.15 68.28 65.29 51.67 53.68 54.74 56.59
@@ 6 Conclusion
The assessment of commonsense accuracy in Section 3.2 relies on automatic labeling by ChatGPT. Therefore, the accuracy presented in Table 1 serves solely for comparing stories and rules, and should not be regarded as absolute accuracy. To ensure the quality of this automatic evaluation, we conduct a manual evaluation of a small set of stories and rules, as shown in Table 9 in Appendix C.1. However, due to its labor-intensive nature, manually labeling all generated stories and rules is impractical for us. Future studies should incorporate more extensive human evaluations to provide a more comprehensive and nuanced understanding of the commonsense knowledge generated by LLMs.
@@
Acknowledgements
We sincerely thank all anonymous reviewers for their insightful comments and valuable suggestions. This research work is supported by the Strategic Priority Research Program of Chinese Academy of Sciences, Grant No. XDA27020200, and the National Natural Science Foundation of China under Grants No. 62122077, 62106251, and 62306303.
@@ Appendix B Details of Commonsense QA Datasets
CommonsenseQA What is another name for a disk for storing information? A. computer store B. computer to store data C. computer hard drive D. cd player E. usb mouse
@@
OpenBookQA Owls spend their nights A. tending to their homes B. sleeping in hollow logs C. scanning their territory for field mice D. hanging out with other owls
@@
PIQA Where can I buy a tennis ball A. You can purchase a tennis ball at any sports store B. You can purchase a tennis racket at any sports store
@@
SocialIQA Aubrey the officer pulled a driver over for speeding on the road. Why did Aubrey do this? A. find a safe place to pull the person over B. so people don’t drive to fast C. look up the person’s license plate number
@@
ARC-easy Scientists at a local university have been studying the impact that people have on Earth. One of the areas being studied is how the burning of fossil fuels affects the environment. Which effect of fossil fuel burning have the scientists most likely evaluated? A. the production of nitrogen-fixing bacteria B. the mechanical weathering of roads C. the formation of acid rain D. the increase in runoff
@@
ARC-challenge How should a line graph be used to display distance and time data for a moving object? A. The y-axis should be labeled as time, which is the dependent variable. B. The y-axis should be labeled as distance, which is the independent variable. C. The x-axis should be labeled as distance, which is the dependent variable. D. The x-axis should be labeled as time, which is the independent variable.
@@
QASC What may renal failure be treated with? A. Laboratory B. Lymphocytes C. saves lives D. dialysis E. Lymph fluid F. dandelions G. ibuprofen H. Protein
@@
AI2Sci-elem To make an electromagnet, a conductor should be coiled around - A. a glass tube B. an iron nail C. a roll of paper D. a wooden stick
@@
AI2Sci-middle Which best describes the characteristics of a river basin? A. the land drained by a river and its tributaries B. the land formed when rivers create estuaries and marshes C. the land at the mouth of a river where water flows into the ocean D. the land formed as a result of a river flooding
@@
WinoGrande She chose the black car over the green car, because the A. black car has more brighter color. B. green car has more brighter color.
@@
WSC The user changed his password from "GrWQWu8JyC" to "willow-towered Canopy Huntertropic wrestles" as A. grwqwu8jyc was easy to remember. B. willow-towered canopy huntertropic wrestles was easy to remember.
NumerSense a french horn has <how many> keys. (three)
@@
HellaSWAG-act.net Another man practices hurling himself backward over a pole onto a gym mat inside of the gym. several more men A. practice hurling street hurling outside and on a gym floor. B. practice hurling while a coach’s hand watches. C. practice long jumps and backward jumps inside of the gym using the sandbox and gym mats as landing tools. D. practice pitches outside of the gym interior.
@@
HellaSWAG-wikihow [header] How to do tiger eye hair [title] Purchase your hair dye. [step] Take a trip to your local drug store or beauty supply store. Depending on the look you’re going for, you may want to keep it simple and just choose one color, or you may want to buy four. A. Your hair will stick out more if you use a thicker dye, such as a mousse or gel. [title] Pour 3 ounces of red wine into your bowl. B. [substeps] In the tattoo artist’s shop you should find several different colors and strips, and apply those to your hair. Make sure the colors match what you want to do. C. It’s totally up to you! You can buy a blonde highlighting kit that will lighten pieces of your brown hair, auburn dye, gorgeous golden hues, soft brown dyes-whichever dye you think will look good in your hair. The darker your hair is, the less of an effect you’ll notice. D. [substeps] If you want to dye your hair yourself, make sure to use the formula before you apply the dye. Gels are sometimes recommended but are highly expensive, and can be difficult to find at grocery stores.
CommonsenseQA2.0 In the US a senator is a person elected to a six year term? A. yes B. no
@@
SWAG Someone finds people playing chess at one of the long polished tables. She A. walks down the brightly decorated hall to join them. B. pats him in the gut with the box. C. faces the building, in wonder, someone and the other recruits stand around watching, uneasy. D. approaches two fat men wearing an earpiece into an office.
@@
Com2Sense Because the drive was 20 miles long, Beth was able to make it to her destination in under 5 minutes. A. True B. False
@@
SciQ What parts of a human possess the highest concentration of thermoreceptors? A. face and hair B. hand and ears C. face and ears D. hands and feet
@@
QuaRel Lebron James a strong player for the Cavs battles Kevin Durant a thin player for a rebound. Who is likely to get the rebound? A. Durant B. Lebron
@@
QuaRTz Long ago the surface of Venus warmed enough that greenhouse gases escaped into the atmosphere. As a result, the greenhouse effect on that planet A. increased B. decreased
@@
CycIC Rob lauded Will. Charity chastised Will. Who made Will feel happy? A. Daisy B. Cliff C. Rob D. Charity E. Joy
@@
ComVE (Task A) Which statement of the two is against common sense? A. The cleaner is in charge of the money at the store B. The cashier is in charge of the money at the store
@@
COPA The woman was in a bad mood. What was the effect of this? A. She engaged in small talk with her friend. B. She told her friend to leave her alone.
@@
PROST A person drops a bottle, a mirror, an egg, and a shirt from a balcony. Which object is the least likely to break? A. bottle B. mirror C. egg D. shirt
CODAH Kieran is a whale. Kieran A. is a mammal B. is a dog C. has six human kids D. is a orange
@@
SCT I wanted to buy a video game console. I asked my parents, and they came up with an idea. They said if I did my chores, I would be given money to save. I did my chores without being asked every week for a whole summer. What is the end of this story? A. My parents gave me enough money to buy the console. B. At the end of the summer I gave the money back to my parents.
@@
$\alpha$NLI The beginning of the story: Jim got ready for his first date.The ending of the story: Since then, she has ignored all of Jim’s text messages. What happened between the begining and the end of the story? A. Jim’s date wasn’t attracted to him. B. Jim went on the date and said he didn’t like the girl.
WinoVenti The walnut was painted. The walnut is A. edible B. toxic
@@
Setting ChatGPT Vicuna Alpaca
Story 96% 94% 88%
Rule 90% 92% 78%
Table 9: Manually evaluated commonsense accuracy of stories and rules.
@@ Appendix C Further Analyses of Generated Stories
C.1 Manual Evaluations of Commonsense Accuracy in Generated Stories and Rules
To further support the commonsense accuracy results presented in Section 3.2, we conduct a manual evaluation of the generated stories and rules. For each model, we sample 50 stories and 50 rules from the Commonsense QA dataset and manually label each one for commonsense correctness. The annotators are the authors of this paper. A story or rule is marked as incorrect if it contains a description that contradicts real-world commonsense. The results, shown in Table 9, indicate that the commonsense accuracy of stories is higher than that of rules across all three models. This further supports our findings in Section 3.2, where ChatGPT is used to automatically assess the commonsense accuracy of a larger set of stories and rules.
C.2 Accuracy Differences between Using Stories and Rules for ChatGPT and Alpaca